If 27200 dollars is invested at an interest rate of 10 percent per year, find the value of the investment at the end of 5 years for the following compounding methods, to the nearest cent.
Annual: $
Semiannual: $
Monthly: $
Daily: $

Answer by Boreal(15235)

the first is 27200(1+0.10/1)^5=$43805.87
the second is 27200(1+(.1/2))^10)$44305.93 Notice how the rate is halved but the. compounding has doubled.
the third is 27200(1+(.1/12))^60=$44752.40
the fourth is 27200(1+(.1/365))^1825=$44842.15
Notice how these slowly increase. There will be a limit at continuously compounding.